OpenStudy (anonymous):

Simplify: [((2x+2h+3)/(x+h)) - ((2x+3)/x)] / h

OpenStudy (jwt625):

=(2+3/(x+h)-2-3/x)/h =(3X-(3x+3h))/((x+h)*x*h) =-3/(x^2+x*h)